# VeFA: Vector-Based Feature Space Adaptation for Mitigating Forgetting in Fine-Tuning

### v3

@@ E.1 Maximum VeFA Scaling Coefficient
Tables 7, 8 and 9 report the maximum VeFA scaling coefficient $\|\boldsymbol{k}\|_{\infty}$ across all layers for each task and backbone. For image classification (Table 7), all datasets yield $\|\boldsymbol{k}\|_{\infty}\leq 1.9$, indicating only mild rescaling of feature dimensions. On GLUE NLU tasks (Table 8), both RoBERTa-Base and RoBERTa-Large exhibit similarly small scaling factors, mostly in the range 1.4–3.9. For NLG with GPT-2 (Table 9), the maximal values remain below 6.2. Overall, these results confirm that VeFA operates with relatively small feature-wise scalings in practice, which is consistent with our theoretical analysis that the forgetting under VeFA is bounded and controlled by $\|\boldsymbol{k}\|_{\infty}$.
@@
E.2 Average Number of Intruder Dimensions
We report the number of intruder dimensions for different GLUE tasks in Fig.5. It can be seen that LoRA consistently introduces intruder dimensions across tasks, which may exacerbate forgetting and cause conflicts when merging adapters, as shown in Tab.3. This also explains why, in our cross-task evaluations, LoRA underperforms VeFA, which by construction does not introduce any additional intruder dimensions. In particular, RTE, STS-B, and MRPC exhibit more intruder dimensions than the other three tasks, likely because they are all fine-tuned from the same MNLI checkpoint.

### v4

@@ F.6 Maximum VeFA Scaling Coefficient
Tables 10, 11 and 12 report the maximum VeFA scaling coefficient $\|\boldsymbol{k}\|_{\infty}$ across all layers for each task and backbone. For image classification (Table 10), all datasets yield $\|\boldsymbol{k}\|_{\infty}\leq 1.9$, indicating only mild rescaling of feature dimensions. On GLUE NLU tasks (Table 11), both RoBERTa-Base and RoBERTa-Large exhibit similarly small scaling factors, mostly in the range 1.4–3.9. For NLG with GPT-2 (Table 12), the maximal values remain below 6.2. Overall, these results confirm that VeFA operates with relatively small feature-wise scalings in practice, which is consistent with our theoretical analysis that the forgetting under VeFA is bounded and controlled by $\|\boldsymbol{k}\|_{\infty}$.
@@
F.7 Analysis of Intruder Dimensions in LoRA
We firstly report the number of intruder dimensions for different GLUE tasks in Fig.8. It can be seen that LoRA consistently introduces intruder dimensions across tasks, which may exacerbate forgetting and cause conflicts when merging adapters, as shown in Tab.9. This also explains why, in our cross-task evaluations, LoRA underperforms VeFA, which introduces no additional intruder dimensions. In particular, RTE, STS-B, and MRPC exhibit more intruder dimensions than the other three tasks, likely because they are all fine-tuned from the same MNLI checkpoint. A layer-wise analysis can be seen in Fig. 9. Intruder directions are present across most layers, with a pronounced concentration in mid-to-late layers for both RTE and MRPC. Moreover, query and value exhibit comparable intruder counts, but their relative dominance varies by layer, indicating that intruder behavior is module- and layer-specific rather than uniformly attributable to a single projection.
Building on the layer-wise statistics of intruder-dimension counts (Fig. 9), we further conduct a finer-grained analysis by decomposing the LoRA update $\boldsymbol{\Delta}$ into two components: (i) the part that lies inside the column space of the pretrained weight $\boldsymbol{W}_{0}$, and (ii) the remaining component residing in the intruder subspace. As shown in Fig. 10, for both RTE and MRPC, the “inside” component consistently dominates the overall update magnitude, while the intruder component is non-negligible and becomes more pronounced in later layers, especially for the query matrices. Notably, the trends in intruder counts and intruder magnitudes are not strictly aligned: a layer may exhibit a moderate number of intruder dimensions yet carry substantial intruder mass (or vice versa), suggesting that “how many” intruder directions appear and “how much” update energy they carry capture complementary aspects of the adaptation behavior.
@@ F.7 Analysis of Intruder Dimensions in LoRA
(a) RTE
(b) MRPC
Figure 9: Number of intruder dimensions in LoRA query and value matrices across layers of RoBERTa-L.
(a) RTE
(b) MRPC
Figure 10: Magnitude of the adapted $\boldsymbol{\Delta}$ vectors in LoRA query and value matrices across layers of RoBERTa-L.